OpenStudy (owen_turner):

What is 4.6 × 108 cm/min in meters per hour? A. 2.76 × 108 m/h B. 2.76 × 1010 m/h C. 7.67 × 106 m/h D. 7.67 × 108 m/h

OpenStudy (triciaal):

you will need 2 conversions cm to m and minutes to hours

OpenStudy (triciaal):

do you know the conversion factors?

OpenStudy (danjs):

Take the given value and units, and multiply it by ratios of 1 to change unit, just remember the way each prefix relates to the base unit, notice the units cancel to leave m/hr \[\large \frac{ 4.6*10^8 cm }{ 1 \min }*\frac{ 10^{-2} m }{ 1 cm }*\frac{ 60 \min }{ 1 hr }\]